OpenStudy (anonymous):

A highway makes an angle of 6 degrees with the horizontal. This angle is maintained for a horizontal distance of 5 miles. To the nearest hundredth of a mile, how high does the highway rise in this 5-mile section? Show the steps you use to find the distance.

OpenStudy (aravindg):

Hi there! Welcome to Openstudy :) If you have any doubts on how the site works ,feel free to ask me or any other Ambassador with purple A near them .Have great time here . Check out our COC : http://openstudy.com/code-of-conduct

OpenStudy (aravindg):

O.K. To your question: Can you try to draw a figure of this problem using the draw button ?

OpenStudy (anonymous):

there is no figure

OpenStudy (aravindg):

I mean can you try to draw a figure using given data ?

OpenStudy (anonymous):

no..... i am not a math person by any means...

OpenStudy (aravindg):

Oh .. Our site aims at getting to the answer by making the asker participate !We just guide you through the process.. ! It is not that hard to draw a figure for this question .Please try

OpenStudy (anonymous):



OpenStudy (anonymous):

i guess i have no idea if this would even be right or not

OpenStudy (aravindg):

Good effort ! Well according to me the fig would be like 

OpenStudy (aravindg):

we need to find h . We can use the trigonometry here . take tan 6 degree for this triangle

OpenStudy (anonymous):

i am in geometry 2 not trigonometry....

OpenStudy (anonymous):

i have no idea what trig is about

OpenStudy (aravindg):

Oh then I think you need to take approximation \[\huge \boxed{\theta =\dfrac{\text{arc length}}{radius}}\]

OpenStudy (aravindg):

here it will be theta =h/5

OpenStudy (aravindg):

first convert 6 degree to radian measure..then use this formula

OpenStudy (anonymous):

ok well thank you for trying to help me but you are speaking a foreign language to me....

OpenStudy (aravindg):

you dont know English ? :/

OpenStudy (anonymous):

wow that was dumb.... yes i know english i dont understand what you are talking about i.e. foriegn language... figure of speech